OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a vertex at the origin and a focus at (0, -4).

OpenStudy (campbell_st):

the focus is below the vertex so the focal length is -4 a form of the parabola is \[x^2 = 4ay\] you know a = -4 so substitute and make y the subject

OpenStudy (anonymous):

y=-1/4x^2?

OpenStudy (campbell_st):

well not quite substitute a = -4 gives \[x^2 = -16 y\] now make y the subject

OpenStudy (jdoe0001):

notice one thing, the vertex is at the origin, that is, (0 , 0) the focus is at (0, -4), lemme graph that  so as you can see, from the vertex to the focus, there's a distance of "4 units" notice the parabola is opening DOWNWARDS, that means a negative focus unit, or -4 and is also moving over the y-axis the equation for a parabola moving over the y-axis using the "focus form" of it will be \(\bf (x-h)^2=4p(y-k)\) (h, k) = vertex coordinates p = distance from the vertex to the focus so plug the values in, and solve for "y" to get the equation

OpenStudy (anonymous):

oh my bad y=-1/16x^2

OpenStudy (campbell_st):

thats it